# Compact Quantum Homogeneous Kähler Spaces

### v4

@@ Abstract.
RÓB is supported by the GAČR/NCN grant Quantum Geometric Representation Theory and Non- commutative Fibrations 24-11728K, and acknowledges support from COST Action 21109 CaLISTA, supported by COST (European Cooperation in Science and Technology) and HORIZON- MSCA-2022-SE-01-01 CaLIGOLA. AC acknowledges support from MSCA-DN CaLiForNIA - 101119552. PS was supported by the GAČR grant GA22-00091S.
@@ 1. Introduction
Since the emergence of quantum groups in the 1980s, a central role in their presentation and development has been played by the theory of operator algebras. We mention in particular Woronowicz’s seminal notion of a compact quantum group [42]. There exists, however, a stark contrast between the development of the noncommutative topological and the noncommutative differential geometric aspects of the theory. For example, for Drinfeld–Jimbo quantum groups, their $C^{*}$-algebraic $K$-theory has long been known to be the same as for their classical counterparts [30]. By contrast, the unbounded formulation of $K$-homology, which is to say Connes and Moscovici’s theory of spectral triples, remains very poorly understood. Indeed, despite a large number of very important contributions over the last thirty years, there is still no consensus on how to construct a spectral triple for $\mathcal{O}_{q}(SU_{2})$, probably the most fundamental example of a quantum group.
@@ 2.2. Hermitian and Kähler Structures
$P^{k}:=\begin{cases}\{\alpha\in\Omega^{k}\,|\,L_{\sigma}^{n-k+1}(\alpha)=0\},&\text{ \leavevmode\nobreak\ if }k\leq n,\\ 0&\text{ \leavevmode\nobreak\ if }k>n.\end{cases}$
@@
We denote the non-zero positive elements of $B$ by $B_{>0}:=B_{\geq 0}\setminus\{0\}$. We say that an Hermitian structure $(\Omega^{(\bullet,\bullet)},\sigma)$ is positive definite if the associated metric $g_{\sigma}$ is positive definite, which is to say, if $g_{\sigma}$ satisfies
@@ 2.3. A Representation of $\mathfrak{sl}_{2}$
$H:\Omega^{\bullet}\to\Omega^{\bullet},$ $\omega\mapsto(k-n)\omega,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{ for }\omega\in\Omega^{k},$
@@ 2.5. Hermitian Modules
$1.\leavevmode\nobreak\ h^{\prime}_{{\mathcal{F}}}(f,k)=h^{\prime}_{{\mathcal{F}}}(k,f)^{*},$ $2.\leavevmode\nobreak\ h^{\prime}_{{\mathcal{F}}}(f,f)\in B_{>0},\textrm{ when }f\neq 0.$
@@
In [33] the pairings $g_{\sigma}$ and $g_{{\mathcal{F}}}$ were chosen to be linear in the first variable, whereas here we take linearity in the second variable. (This is required in order that the representation in Corollary 3.10 is a *-map.) However, as a careful reading of the [33] will confirm, this does not affect any of the results used above.
@@ 3.3. Morphisms as Bounded Operators
Since $L_{\sigma}\otimes\mathrm{id}_{{\mathcal{F}}},\leavevmode\nobreak\ \Lambda_{\sigma}\otimes\mathrm{id}_{{\mathcal{F}}}$, and $H\otimes\mathrm{id}_{{\mathcal{F}}}$ are all morphisms in $^{A}_{B}\mathrm{mod}_{0}$, Proposition 3.4 implies that they extend to bounded operators on $L^{2}(\Omega^{\bullet}\otimes_{B}{\mathcal{F}})$. It now follows from the $\mathfrak{sl}_{2}$-representation given in [32, Corollary 5.14] that we get a bounded Lie algebra representation of $\mathfrak{sl}_{2}$. ∎
@@ 3.4. Bounded Multiplication Maps
With this result in hand, we now show that $L_{\omega}$ is a bounded linear operator. For any element $\nu\otimes f$ of $\Omega^{\bullet}\otimes_{B}{\mathcal{F}}$,
$\mathrm{U}\circ L_{\omega}\circ\mathrm{U}^{-1}\left(\nu_{(-1)}f_{(-1)}\otimes[\nu_{(0)}]\otimes[f_{(0)}]\right)=$ $\,\mathrm{U}\circ L_{\omega}(\nu\otimes f)$
$=$ $\,\mathrm{U}(\omega\wedge\nu\otimes f)$
@@
$=$ $\,(\lambda_{A}(\omega_{(-1)})\otimes l_{\omega_{(0)}})(\nu_{(-1)}f_{(-1)}\otimes[\nu_{(0)}]\otimes[f_{(0)}]).$
@@
Finally, we observe that since $\Omega^{\bullet}\otimes_{B}{\mathcal{F}}$ is dense in $L^{2}(\Omega^{\bullet}\otimes_{B}{\mathcal{F}})$ by construction, $L_{\omega}$ uniquely extends to an element of $\mathbb{B}(L^{2}(\Omega^{\bullet}\otimes_{B}{\mathcal{F}})$. ∎
@@ 4.3. Cores and Domains
Since we are assuming that $A$ is finitely generated as an algebra, it follows from Proposition 3.2 that $L^{2}(\Omega^{\bullet}\otimes_{B}{\mathcal{F}})$ is separable. The subspace $\Omega^{\bullet}\otimes_{B}{\mathcal{F}}\subseteq\mathrm{dom}(D_{\overline{\partial}_{\mathcal{F}}})$ is a core by construction of the closure of $D_{\overline{\partial}_{{\mathcal{F}}}}$. The core is clearly closed under the action of $\lambda(b)$, for all $b\in B$. Corollary 3.12 says that $D_{\overline{\partial}_{{\mathcal{F}}}},\lambda(b)]$ is a bounded operator on $\Omega^{\bullet}\otimes_{B}{\mathcal{F}}$, for all $b\in B$, and so, we see that $\lambda(b)\mathrm{dom}(D_{\overline{\partial}_{{\mathcal{F}}}})$ is contained in $\mathrm{dom}(D_{\overline{\partial}_{{\mathcal{F}}}})$ as claimed. ∎
@@ 5.1. The Holomorphic Euler Characteristic
where we have denoted by $H_{\overline{\partial}_{{\mathcal{F}}}}^{(0,k)}$ the $k$-cohomology group of the complex of twisted anti-holomorphic forms $\overline{\partial}_{{\mathcal{F}}}:\Omega^{(0,\bullet)}\otimes_{B}{\mathcal{F}}\to\Omega^{(0,\bullet)}\otimes_{B}{\mathcal{F}}$.
@@ 6.3. Relative Line Modules over the Irreducible Quantum Flag Manifolds
Table 1. Irreducible Quantum Flag Manifolds: organised by series, with defining crossed node numbered according to [23, §11.4], CQGA-homogeneous space symbol and name
@@
gives ${\mathcal{E}}_{l}$ the structure of a covariant Hermitian relative line module. Since ${\mathcal{E}}_{l}$ is a simple object in $^{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{O}_{q}(G)}_{\mathcal{O}_{q}(G/L_{S})}\textrm{mod}_{0}$, we see that $h_{{\mathcal{E}}_{l}}$ is the unique such structure up to positive scalar multiple. Finally, we recall from [16, Theorem 4.9] that for $k>0$, the line module ${\mathcal{E}}_{k}$ is positive with respect to $\kappa$, and ${\mathcal{E}}_{-k}$ is negative.
@@ 6.4. Compact Quantum Homogeneous Kähler Spaces
$\Omega^{\bullet}_{q}(G/L_{S})\in\,{}^{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{O}_{q}(G)}_{\mathcal{O}_{q}(G/L_{S})}\textrm{mod}_{0}$
@@ 6.6. Twisted Dolbeault–Dirac Fredholm Operators
We denote the associated Chern connection by $\nabla$, and note that it is a bimodule connection [6, §6.5]. It now follows from the results of §3.5 that the commutators $\nabla,\lambda(b)]$,   $\partial_{{\mathcal{F}}},\lambda(b)$, and $\overline{\partial}_{{\mathcal{F}}},\lambda(b)$ are bounded operators. Moreover, it follows from the results of §4.2 that the operators $\nabla,\partial_{{\mathcal{F}}}$, and $\overline{\partial}_{{\mathcal{F}}}$ are essentially self-adjoint, as are the corresponding Laplacians.
@@
(9) $\mathrm{index}\!\left(D_{\overline{\partial}_{{\mathcal{E}}_{-k}}}\right)=\begin{cases}\leavevmode\nobreak\ (-1)^{M}\dim\!\big{(}V_{(k-C_{S})\varpi_{x}}\big{)},&\textrm{ if }k\geq C_{S},\\ \leavevmode\nobreak\ 0,&\textrm{ otherwise},\end{cases}$
@@
$H^{(0,b)}_{\overline{\partial}_{{\mathcal{E}}_{-k}}}=0,$  $\textrm{ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for all }b<M.$
Moreover, it was shown in [33] that Serre duality, together with the Borel–Weil theorem for the irreducible quantum flag manifolds [9, Theorem 6.1], implies that cohomology group $H^{(0,M)}_{\overline{\partial}_{{\mathcal{E}}_{-k}}}$ is a finite-dimensional irreducible $U_{q}(\mathfrak{g})$-module. If $k\geq C_{S}$, then it has highest weight $w_{0}((C_{S}-k)\varpi_{x})$, and otherwise it is the zero module. In particular, the anti-holomorphic cohomology groups of ${\mathcal{E}}_{-k}$ are finite-dimensional, implying that the ${\mathcal{E}}_{-k}$-twisted Dolbeault–Dirac operator is a Fredholm operator as claimed.
@@ A.1. Compact Quantum Group Algebras
$\mathcal{C}(V):=\mathrm{span}_{\mathbb{C}}\{(\mathrm{id}\otimes f)\Delta_{L}(v)\,|\,f\in\text{Hom}_{\mathbb{C}}(V,\mathbb{C}),v\in V\}\subseteq A.$
@@ C.1. Spectral Triples and the Bounded Transform
One reason why spectral triples are important is because they provide unbounded representatives for $K$-homology classes. For a spectral triple $(A,\mathcal{H},D)$, its bounded transform is the operator
@@ C.2. Spectral Triples and Dolbeault–Dirac Eigenvalues
We now formulate precise criteria for when the Dolbeault–Dirac operator of a CQH-Hermitian space gives a spectral triple. For sake of clarity and convenience, let us recall the relevant properties of $L^{2}(\Omega^{\bullet})$ and $D_{\overline{\partial}}$. If $A$ is finitely generated, then it follows from Proposition 3.2 that $L^{2}(\Omega^{\bullet})$ is separable. By Corollary 3.13 we have a faithful *-representation $\lambda:B\to\mathbb{B}(L^{2}(\Omega^{\bullet}))$. From Corollary 4.6 we know that $D_{\overline{\partial}}$ is an essentially self-adjoint operator, which is, moreover, densely-defined by construction. By Corollary 3.12, the commutators $D_{\overline{\partial}},\lambda(b)]$ are bounded, and by Proposition 4.8 above, $\lambda(b)\mathrm{dom}(D_{\overline{\partial}})\subseteq\mathrm{dom}(D_{\overline{\partial}})$, for all $b\in B$. With respect to the obvious $\mathbb{Z}_{2}$-grading $\gamma$, the operator $D_{\overline{\partial}}$ is of degree $1$, and $\lambda(b)$ is a degree 0 operator, for all $b\in B$. Finally, we note that since $D_{\overline{\partial}}$ is diagonalisable on $L^{2}(\Omega^{\bullet})$, it has compact resolvent if and only if its eigenvalues tend to infinity and have finite multiplicity. Collecting these facts together gives the following proposition.

### v5

@@ Abstract.
RÓB is supported by the GAČR/NCN grant Quantum Geometric Representation Theory and Non- commutative Fibrations 24-11728K, and acknowledges support from COST Action 21109 CaLISTA, supported by COST (European Cooperation in Science and Technology) and HORIZON- MSCA-2022-SE-01-01 CaLIGOLA. PS was supported by the GAČR grant GA22-00091S.
@@ 1. Introduction
Since the emergence of quantum groups in the 1980s, a central role in their presentation and development has been played by the theory of operator algebras. We mention in particular Woronowicz’s seminal notion of a compact quantum group [42, 43]. There exists, however, a stark contrast between the development of the noncommutative topological and the noncommutative differential geometric aspects of the theory. For example, for Drinfeld–Jimbo quantum groups, their $C^{*}$-algebraic $K$-theory has long been known to be the same as for their classical counterparts [30]. By contrast, the unbounded formulation of $K$-homology, which is to say Connes and Moscovici’s theory of spectral triples, remains very poorly understood. Indeed, despite a large number of very important contributions over the last thirty years, there is still no consensus on how to construct a spectral triple for $\mathcal{O}_{q}(SU_{2})$, probably the most fundamental example of a quantum group.
@@ 2.2. Hermitian and Kähler Structures
$P^{k}:=\begin{cases}\{\alpha\in\Omega^{k}\,|\,L_{\sigma}^{n-k+1}(\alpha)=0\},&\text{ \penalty 10000\ if }k\leq n,\\ 0&\text{ \penalty 10000\ if }k>n.\end{cases}$
@@
[…]
@@ 2.3. A Representation of $\mathfrak{sl}_{2}$
$H:\Omega^{\bullet}\to\Omega^{\bullet},$ $\omega\mapsto(k-n)\omega,\penalty 10000\ \penalty 10000\ \text{ for }\omega\in\Omega^{k},$
@@ 2.5. Hermitian Modules
$1.\penalty 10000\ h^{\prime}_{{\mathcal{F}}}(f,k)=h^{\prime}_{{\mathcal{F}}}(k,f)^{*},$ $2.\penalty 10000\ h^{\prime}_{{\mathcal{F}}}(f,f)\in B_{>0},\textrm{ when }f\neq 0.$
@@
In [33] the pairings $g_{\sigma}$ and $g_{{\mathcal{F}}}$ were chosen to be linear in the first variable, whereas here we take linearity in the second variable. (This is required in order that the representation in Proposition 3.10 is a *-map.) However, as a careful reading of the [33] will confirm, this does not affect any of the results used above.
@@ 3.3. Morphisms as Bounded Operators
Since $L_{\sigma}\otimes\mathrm{id}_{{\mathcal{F}}},\penalty 10000\ \Lambda_{\sigma}\otimes\mathrm{id}_{{\mathcal{F}}}$, and $H\otimes\mathrm{id}_{{\mathcal{F}}}$ are all morphisms in $^{A}_{B}\mathrm{mod}_{0}$, Proposition 3.4 implies that they extend to bounded operators on $L^{2}(\Omega^{\bullet}\otimes_{B}{\mathcal{F}})$. It now follows from the $\mathfrak{sl}_{2}$-representation given in [32, Corollary 5.14] that we get a bounded Lie algebra representation of $\mathfrak{sl}_{2}$. ∎
@@ 3.4. Bounded Multiplication Maps
With this result in hand, we now show that $L_{\omega}$ is a bounded linear operator. For any element $\nu\otimes f$ of $\Omega^{\bullet}\otimes_{B}{\mathcal{F}}$, we see that
$\mathrm{U}\circ L_{\omega}\circ\mathrm{U}^{-1}\left(\nu_{(-1)}f_{(-1)}\otimes[\nu_{(0)}]\otimes[f_{(0)}]\right)$
simplifies to the expression
$\mathrm{U}\circ L_{\omega}(\nu\otimes f)=$ $\,\mathrm{U}(\omega\wedge\nu\otimes f)$
@@
$=$ $\,((\lambda_{A}(\omega_{(-1)})\otimes l_{\omega_{(0)}})(\nu_{(-1)}f_{(-1)}\otimes[\nu_{(0)}]\otimes[f_{(0)}]).$
@@
Finally, we observe that since $\Omega^{\bullet}\otimes_{B}{\mathcal{F}}$ is dense in $L^{2}(\Omega^{\bullet}\otimes_{B}{\mathcal{F}})$ by construction, $L_{\omega}$ uniquely extends to an element of $\mathbb{B}(L^{2}(\Omega^{\bullet}\otimes_{B}{\mathcal{F}}))$. ∎
@@ 4.3. Cores and Domains
Since we are assuming that $A$ is finitely generated as an algebra, it follows from Proposition 3.2 that $L^{2}(\Omega^{\bullet}\otimes_{B}{\mathcal{F}})$ is separable. The subspace $\Omega^{\bullet}\otimes_{B}{\mathcal{F}}\subseteq\mathrm{dom}(D_{\overline{\partial}_{\mathcal{F}}})$ is a core by construction of the closure of $D_{\overline{\partial}_{{\mathcal{F}}}}$. The core is clearly closed under the action of $\lambda(b)$, for all $b\in B$. Proposition 3.12 says that $D_{\overline{\partial}_{{\mathcal{F}}}},\lambda(b)]$ is a bounded operator on $\Omega^{\bullet}\otimes_{B}{\mathcal{F}}$, for all $b\in B$, and so, we see that $\lambda(b)\mathrm{dom}(D_{\overline{\partial}_{{\mathcal{F}}}})$ is contained in $\mathrm{dom}(D_{\overline{\partial}_{{\mathcal{F}}}})$ as claimed. ∎
@@ 5.1. The Holomorphic Euler Characteristic
where we have denoted by $H_{\overline{\partial}_{{\mathcal{F}}}}^{(0,k)}$ the $k$-cohomology group of the complex of twisted anti-holomorphic forms
$\overline{\partial}_{{\mathcal{F}}}:\Omega^{(0,\bullet)}\otimes_{B}{\mathcal{F}}\to\Omega^{(0,\bullet)}\otimes_{B}{\mathcal{F}}.$
@@ 6.3. Relative Line Modules over the Irreducible Quantum Flag Manifolds
Table 1. Irreducible Quantum Flag Manifolds: organised by series, with defining crossed node, CQGA-homogeneous space symbol, and name.
@@
For the reader’s convenience, we recall in the above table the standard pictorial description of the quantum Levi subalgebras defining the irreducible quantum flag manifolds, given in terms of Dynkin diagrams.
@@
gives ${\mathcal{E}}_{l}$ the structure of a covariant Hermitian relative line module. Since ${\mathcal{E}}_{l}$ is a simple object in the abelian category
$^{\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{O}_{q}(G)}_{\mathcal{O}_{q}(G/L_{S})}\textrm{mod}_{0}$,
we see that $h_{{\mathcal{E}}_{l}}$ is the unique such structure up to positive scalar multiple. Finally, we recall from [16, Theorem 4.9] that for $k>0$, the line module ${\mathcal{E}}_{k}$ is positive with respect to $\kappa$, and ${\mathcal{E}}_{-k}$ is negative.
@@ 6.4. Compact Quantum Homogeneous Kähler Spaces
$\Omega^{\bullet}_{q}(G/L_{S})\in\,{}^{\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{O}_{q}(G)}_{\mathcal{O}_{q}(G/L_{S})}\textrm{mod}_{0}$
@@ 6.6. Twisted Dolbeault–Dirac Fredholm Operators
We denote the associated Chern connection by $\nabla$, and note that it is a bimodule connection [6, §6.5]. It now follows from the results of §3.5 that the commutators $\nabla,\lambda(b)]$,   $\partial_{{\mathcal{F}}},\lambda(b)]$, and $\overline{\partial}_{{\mathcal{F}}},\lambda(b)]$ are bounded operators. Moreover, it follows from the results of §4.2 that the operators $\nabla,\partial_{{\mathcal{F}}}$, and $\overline{\partial}_{{\mathcal{F}}}$ are essentially self-adjoint, as are the corresponding Laplacians.
@@
(9) $\mathrm{index}\!\left(D_{\overline{\partial}_{{\mathcal{E}}_{-k}}}\right)=\begin{cases}\penalty 10000\ (-1)^{M}\dim\!\big(V_{(k-C_{S})\varpi_{x}}\big),&\textrm{ if }k\geq C_{S},\\ \penalty 10000\ 0,&\textrm{ otherwise},\end{cases}$
@@
$H^{(0,b)}_{\overline{\partial}_{{\mathcal{E}}_{-k}}}=0,$ $\textrm{ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ for all }b<M.$
Moreover, it was shown in [33] that Serre duality, together with the Borel–Weil theorem for the irreducible quantum flag manifolds [9, Theorem 6.1], implies that
$H^{(0,M)}_{\overline{\partial}_{{\mathcal{E}}_{-k}}}$
is a finite-dimensional irreducible $U_{q}(\mathfrak{g})$-module. If $k\geq C_{S}$, then it has highest weight $w_{0}((C_{S}-k)\varpi_{x})$, and otherwise it is the zero module. In particular, the anti-holomorphic cohomology groups of ${\mathcal{E}}_{-k}$ are finite-dimensional, implying that the ${\mathcal{E}}_{-k}$-twisted Dolbeault–Dirac operator is a Fredholm operator as claimed.
@@
$H^{(0,n)}_{\overline{\partial}_{{\mathcal{E}}_{-k}}}$
of the relative line module ${\mathcal{E}}_{-k}$ vanishes, meaning that the index of its twisted Dirac operator is zero. For the line module ${\mathcal{E}}_{-n-1}$, the dimension of
$H^{(0,n)}_{\overline{\partial}_{{\mathcal{E}}_{-n-1}}}$
is $1$, meaning that its twisted Dirac operator has index $(-1)^{n}$. For the line module ${\mathcal{E}}_{-n-2}$, the top cohomology group
$H^{(0,n)}_{\overline{\partial}_{{\mathcal{E}}_{-n-2}}}$
has the dimension of the representation $V_{\varpi_{1}}$, which is to say, the dimension of the vector space representation of $\mathfrak{sl}_{n+1}$. Thus we see that
@@
$H^{(0,s(n+1-s))}_{\overline{\partial}_{{\mathcal{E}}_{-k}}}$,
meaning that the index of the twisted Dirac operator is zero. For the line module ${\mathcal{E}}_{-n-1}$, the dimension of
$H^{(0,s(n+1-s))}_{\overline{\partial}_{{\mathcal{E}}_{-n}}}$
is $1$, meaning that its twisted Dirac operator has index $(-1)^{s(n+1-s)}$. For the relative line module ${\mathcal{E}}_{-n-2}$, the top anti-holomorphic cohomology group has the same dimension as the fundamental representation $V_{\varpi_{s}}$, which is to say, the dimension of the $s$-fold exterior power of the vector space representation of $\mathfrak{sl}_{n+1}$. Thus we see that
@@
$H^{(0,\frac{n(n-1)}{2})}_{\overline{\partial}_{{\mathcal{E}}_{-k}}}$
vanishes, meaning that the index of the associated twisted Dirac operator is zero. Taking the line module ${\mathcal{E}}_{-2(n-1)}$, the cohomology group has dimension $1$, meaning that its twisted Dirac operator has index $(-1)^{\frac{n(n-1)}{2}}$. For the line module ${\mathcal{E}}_{-2n+1}$, the cohomology group has the same dimension as the representation $V_{\varpi_{n}}$, one of the two half-spin representations. Thus we have that
@@ A.1. Compact Quantum Group Algebras
$\mathcal{C}(V):=\mathrm{span}_{\mathbb{C}}\Big\{(\mathrm{id}\otimes f)\Delta_{L}(v)\,|\,f\in\text{Hom}_{\mathbb{C}}(V,\mathbb{C}),v\in V\Big\}\subseteq A.$
@@ C.1. Spectral Triples and the Bounded Transform
One reason why spectral triples are important is that they provide unbounded representatives for $K$-homology classes. For a spectral triple $(A,\mathcal{H},D)$, its bounded transform is the operator
@@ C.2. Spectral Triples and Dolbeault–Dirac Eigenvalues
We now formulate precise criteria for when the Dolbeault–Dirac operator of a CQH-Hermitian space gives a spectral triple. For sake of clarity and convenience, let us recall the relevant properties of $L^{2}(\Omega^{\bullet})$ and $D_{\overline{\partial}}$. If $A$ is finitely generated, then it follows from Proposition 3.2 that $L^{2}(\Omega^{\bullet})$ is separable. By Proposition 3.10 we have a faithful *-representation $\lambda:B\to\mathbb{B}(L^{2}(\Omega^{\bullet}))$. From Corollary 4.6 we know that $D_{\overline{\partial}}$ is an essentially self-adjoint operator, which is, moreover, densely-defined by construction. By Proposition 3.12, the commutators $D_{\overline{\partial}},\lambda(b)]$ are bounded, and by Proposition 4.8 above, $\lambda(b)\mathrm{dom}(D_{\overline{\partial}})\subseteq\mathrm{dom}(D_{\overline{\partial}})$, for all $b\in B$. With respect to the obvious $\mathbb{Z}_{2}$-grading $\gamma$, the operator $D_{\overline{\partial}}$ is of degree $1$, and $\lambda(b)$ is a degree 0 operator, for all $b\in B$. Finally, we note that since $D_{\overline{\partial}}$ is diagonalisable on $L^{2}(\Omega^{\bullet})$, it has compact resolvent if and only if its eigenvalues tend to infinity and have finite multiplicity. Collecting these facts together gives the following proposition.